OpenStudy (anonymous):

Use Gaussian elimination to find the complete solution to the system of equations, or state that none exists. 4x - y + 3z = 12 x + 4y + 6z = -32 5x + 3y + 9z = 20 A. {(8, -7, -2)} B. {(-8, -7, 9)} C. ∅ D. {(2, -7, -1)}

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I recommend using the linear algebra toolkit provided in the link below http://www.math.odu.edu/~bogacki/lat/

jimthompson5910 (jim_thompson5910):

that will perform Gaussian elimination and will do so step by step

OpenStudy (anonymous):

UM OK

jimthompson5910 (jim_thompson5910):

let me know if you need help using it

OpenStudy (anonymous):

YEAH I DO..I FORGET THINGS EASILY

jimthompson5910 (jim_thompson5910):

ok click the "Enter" link

jimthompson5910 (jim_thompson5910):

then click "Row operation calculator"

OpenStudy (anonymous):

I DID

jimthompson5910 (jim_thompson5910):

pick 3 rows then 4 columns

OpenStudy (anonymous):

OK DONE

jimthompson5910 (jim_thompson5910):

now you should see a blank 3x4 matrix enter the data values in the boxes

OpenStudy (anonymous):

ENTER THESE 4x - y + 3z = 12 x + 4y + 6z = -32 5x + 3y + 9z = 20

jimthompson5910 (jim_thompson5910):

yes, but you do NOT enter the variables

jimthompson5910 (jim_thompson5910):

so in the first row, you would enter 4, -1, 3, 12 one number per box (in the first row)

jimthompson5910 (jim_thompson5910):

the other rows are entered in a similar way

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

SO THE SECOND WOULD BE

OpenStudy (anonymous):

1 4 6 -32

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

one number per box

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

I SUBMITTED IT

jimthompson5910 (jim_thompson5910):

now click on the button that says "Reduced Row Echelon Form"

jimthompson5910 (jim_thompson5910):

it's the button on the far right

OpenStudy (anonymous):

OK I DID

jimthompson5910 (jim_thompson5910):

what you'll get are the full steps if you performed Gaussian elimination

jimthompson5910 (jim_thompson5910):

notice on the very last step (the matrix on the right) we have this row 3 0 0 0 1 this implies that 0x + 0y + 0z = 1 which turns into 0 = 1 but that's FALSE, so there are NO solutions and this system is inconsistent

jimthompson5910 (jim_thompson5910):

to represent the idea of "no solutions", we refer to the empty set { } or ∅

OpenStudy (anonymous):

I UNDERSTAND WHAT YOU ARE SAYING NOW THANK YOU FOR ALL YOUR HELP @jim_thompson5910 ;-)

jimthompson5910 (jim_thompson5910):

you're welcome